# Retrieval-Augmented Generation; Entity Linking; Domain Adaptation; Wikidata; Educational Question Answering

### v1

@@ 4.2 Method 2: Subjective Evaluation through LLM-based Scoring
Analysis of Response Quality. Consistent with the retrieval metrics observed in Method 1, the ELERAG (RRF) strategy achieves the highest scores across all qualitative dimensions. The improvement is particularly noticeable in Completeness ($6.10$) and Relevance ($5.57$). This indicates that the chunks prioritized by the RRF algorithm—by balancing semantic similarity and entity popularity—provide the generator with a richer and more pertinent context, allowing it to construct answers that are not only factually correct but also more exhaustive.
Comparison with Cross-Encoder. Interestingly, the RRF+Cross-Encoder pipeline scores slightly lower than both ELERAG and the Baseline in terms of qualitative scoring. This reinforces the hypothesis derived from the quantitative analysis: although the Cross-Encoder is a powerful semantic filter, in this specific educational domain it may be overly aggressive, discarding chunks that contain necessary context or entity definitions that are valuable for the final generation. Consequently, the simpler and more robust RRF fusion proves to be superior for end-to-end answer quality.
Clarity Stability. The scores for Clarity are relatively stable across all systems (ranging from $4.42$ to $4.54$). Since the evaluation target in this method is formed by directly concatenating the raw retrieved chunks, this stability is expected: the source material (lecture transcripts) is identical for all pipelines. However, the slight edge held by ELERAG ($4.54$) suggests that the entity-driven retrieval tends to select more structurally complete or self-contained passages, resulting in a slightly more coherent reading flow compared to the other methods.
@@ 4.3 Method 3: Classical Evaluation on the Full RAG Pipeline
Impact of Entity-Aware Ranking on Generation. The results confirm that the retrieval quality improvement observed in the dense stage propagates effectively to the final RAG output. ELERAG (RRF) consistently outperforms all other configurations, achieving the highest scores in Exact Match ($0.551$), Recall ($0.589$), and MRR ($0.622$). This indicates that when the LLM is fed with chunks prioritized by our RRF-Based strategy, it is more likely to identify, use, and cite the correct gold answer.
Performance Drop of Cross-Encoder. A critical observation is the underperformance of the RRF+Cross-Encoder pipeline, which records the lowest Exact Match ($0.507$) and MRR scores among all entity-enhanced methods, falling even slightly behind the Baseline. Even with the LLM acting as a final filter, the Cross-Encoder fails to recover the gap. This reinforces the conclusion that for specialized educational content, a heavy semantic re-ranker may introduce noise or drift away from the specific factual precision required, whereas the explicit entity signal used in ELERAG ensures a more robust alignment with the query’s intent.
4.4 Method 1: Evaluation on the Standard Benchmark (SQuAD-it)
Finally, Table 4 reports the retrieval performance on the SQuAD-it dataset. This benchmark serves as a control experiment on a general-domain corpus (Wikipedia).
@@ 4.4 Method 1: Evaluation on the Standard Benchmark (SQuAD-it)
Cross-Encoder 0.777 0.777 0.885 0.912 0.936 0.777 0.836
Analysis of Results.
On the SQuAD-it dataset, the trend observed in the educational benchmark is sharply reversed. The RRF + Cross-Encoder pipeline achieves the best performance across all metrics, with an Exact Match of $0.777$ and an MRR of $0.836$, significantly outperforming both the Baseline and the entity-based methods. Conversely, the ELERAG (RRF) method performs slightly worse than the Baseline ($0.672$ vs $0.693$ in EM).
Evidence of Domain Mismatch. This divergence in performance between the two datasets provides strong experimental evidence for the Domain Mismatch hypothesis. The Cross-Encoder model (mMiniLM) was pre-trained on massive general-domain datasets (MS MARCO, Wikipedia), which are linguistically very similar to SQuAD-it. Consequently, it can leverage its internal knowledge to rank these documents effectively. However, as shown in Section 4, this advantage disappears when applied to the specialized, high-ambiguity domain of university lectures, where our proposed ELERAG method prevails.
@@
This leads to a crucial conclusion: while Cross-Encoders are the State-of-the-Art solution for general web-search tasks, they are not universally optimal. For specialized educational platforms, where training data for fine-tuning is scarce, our entity-aware hybrid approach provides a more robust and accurate retrieval mechanism without the computational cost of a heavy neural re-ranker.
5 Discussion
5.1 Interpretation of Results
@@
The experimental results present a clear and divergent pattern across the two benchmarks, which forms the central finding of this study.
First, the quantitative analysis on our specialized educational corpus demonstrates that our proposed ELERAG architecture achieved the best overall performance. It consistently outperformed not only the Weighted-Score baseline but also the SOTA Cross-Encoder model in key retrieval metrics, such as Exact Match and Mean Reciprocal Rank (MRR). Conversely, on the general-purpose SQuAD-it benchmark, this trend inverted sharply: the Cross-Encoder configuration obtained the highest scores across all metrics, proving its superiority in a standard, general-domain QA setting.
@@
This quantitative divergence provides strong experimental evidence for our Domain Mismatch hypothesis. It indicates that while large, pre-trained re-rankers excel on web-style data (like Wikipedia), a domain-adapted hybrid model like ELERAG is significantly more effective on specialized, narrative corpora (like university lectures). In these contexts, the explicit signals provided by Entity Linking align better with the data than the generic semantic patterns learned by the Cross-Encoder.
@@
Having established that ELERAG is the most effective solution for the specific domain, qualitative analysis helps to explain why the integration of Entity Linking provides this advantage. Inspection of the results confirms that the entity-based signal is crucial in high-ambiguity scenarios. A clear example is the query “Who is Smith?”. The Baseline system, relying only on semantic similarity, retrieved scattered and less relevant documents based on broad keyword matching. In contrast, the ELERAG pipeline, guided by the unambiguous entity ID from Wikidata, successfully promoted the correct, relevant passages to the top ranks. This allowed the LLM to generate a more coherent, factually dense, and precise answer. This qualitative benefit was less pronounced in broad, low-ambiguity queries, where the baseline’s semantic search was already sufficient. This suggests that the primary value of our hybrid approach lies in its ability to resolve factual ambiguity, which is a critical weakness of purely semantic systems in technical domains.
@@
5.2 Implications and Contributions
The results offer several important implications for research in retrieval-augmented generation and educational AI. First, the consistent improvements obtained through entity-based re-ranking demonstrate that factual grounding can be enhanced without the need for expensive model retraining, solely through structured post-retrieval refinement. Second, the contrast between ELERAG and Cross-Encoder performance emphasizes that high-capacity models do not generalize uniformly across all domains: their success strongly depends on the alignment between pre-training data and the target corpus. This finding suggests that lightweight hybrid methods—such as our entity-enriched retrieval—can be a more efficient and accurate alternative to SOTA re-rankers in domain-specific applications, particularly in low-resource or multilingual settings such as Italian educational materials. Finally, the integration of an Entity Linking module provides a clear framework for combining symbolic and neural representations, showing how explicit knowledge bases like Wikidata can complement dense embeddings to improve interpretability and precision.
5.3 Limitations and Future Work
Despite the promising results, this study has limitations that open clear avenues for future research. A primary limitation is the size and source of our custom benchmark; while effective for this study, its automatic generation via LLM APIs (even with validation) may introduce inherent biases compared to human-curated datasets. Furthermore, our analysis did not include a systematic evaluation of computational latency, although the architectural simplicity of RRF compared to Cross-Encoders suggests a theoretical advantage.
These limitations directly inform future work. The most critical next step, which follows directly from our “Domain Mismatch” finding, is to fine-tune a Cross-Encoder model specifically on our educational corpus. We hypothesize that this domain-adaptation step would allow the Cross-Encoder to learn the specific linguistic patterns of the lectures, potentially surpassing the RRF system and combining the best of both worlds: domain knowledge and SOTA neural architecture.
Further research could also explore adaptive weighting schemes for the RRF, dynamically adjusting the balance between semantic and entity signals based on the query type (e.g., factual vs. conceptual). Finally, expanding the evaluation to multilingual data and incorporating human-in-the-loop assessments would provide a richer understanding of the system’s real-world educational value.
@@ Acknowledgements.
GPT-40 Generative Pre-trained Transformer 4.0

### v2

@@ 4.2 Method 2: Subjective Evaluation through LLM-based Scoring
Cross-Encoder 5.83 5.30 4.32
Consistent with the quantitative metrics, the ELERAG (RRF) strategy achieved the highest scores across all qualitative dimensions. The improvement is particularly substantial in Completeness ($6.10$) and Relevance ($5.57$). This indicates that balancing semantic similarity with entity popularity provides the generator with a richer and more pertinent context, allowing it to construct answers that are not only factually correct but also exhaustive.
The qualitative evaluation also highlights the limitations of general-purpose neural re-rankers in this specialized domain. The Standalone Cross-Encoder recorded the lowest scores for Completeness and Relevance. This suggests that while the Cross-Encoder identifies semantically related content, it prioritizes chunks that are tangentially relevant but lack the specific factual definitions required for educational quality—effectively introducing "semantic noise" that the robust RRF fusion successfully filters out.
The scores for Clarity are relatively stable across all systems (ranging from $4.42$ to $4.54$). Since the evaluation target in this method is formed by directly concatenating the raw retrieved chunks, this stability is expected: the source material (lecture transcripts) is identical for all pipelines. However, the slight edge held by ELERAG ($4.54$) suggests that the entity-driven retrieval tends to select more structurally complete or self-contained passages, resulting in a slightly more coherent reading flow compared to the other methods.
@@ 4.3 Method 3: Classical Evaluation on the Full RAG Pipeline
Cross-Encoder 0.522 0.589 0.437 0.599 0.713
The results confirm that the retrieval quality improvements propagate effectively to the final RAG output. ELERAG (RRF) consistently outperforms all other configurations, achieving the highest scores in Exact Match ($0.551$) and MRR ($0.622$). Notably, while it ties with the Standalone Cross-Encoder for the highest Recall ($0.589$), it converts this into significantly better Precision ($0.458$ vs $0.437$). This demonstrates that the Entity-Aware RRF strategy provides a cleaner context, enabling the LLM to identify and cite the correct gold answer more frequently.
The comparison with the Standalone Cross-Encoder further reinforces the importance of ranking precision. Despite matching ELERAG in retrieval recall, the Cross-Encoder performs identically to the Baseline in Exact Match ($0.522$). This discrepancy suggests that the neural model successfully retrieves relevant documents but fails to rank the specific gold chunk effectively, leading the LLM to synthesize information from a mix of relevant documents rather than focusing on the single best source. The RRF+Cross-Encoder pipeline showed a slight degradation (EM $0.507$), confirming that increasing complexity does not yield additive benefits in this setup.
4.4 Evaluation on a standard benchmark (SQuAD-it)
Table 4 reports the retrieval performance on the SQuAD-it dataset. This benchmark serves as a control experiment on a general-domain corpus (Wikipedia).
@@ 4.4 Evaluation on a standard benchmark (SQuAD-it)
RRF+Cross-Encoder 0.777 0.777 0.885 0.912 0.936 0.777 0.836
Cross-Encoder 0.776 0.776 0.885 0.912 0.938 0.776 0.836
On the SQuAD-it dataset, the performance trend is sharply reversed. The Cross-Encoder configurations achieved the best results across all metrics (EM $\approx 0.777$, MRR $0.836$), significantly outperforming both the baseline and the entity-based methods (ELERAG EM $0.672$). This can be explained by the fact that this dataset represents a highly in-domain scenario. SQuAD-it is derived from Wikipedia, which closely resembles the type of data commonly used to train or pre-train many embedding models and cross-encoders. In such settings dense and cross-encoder models tend to perform particularly well because the lexical, stylistic, and semantic characteristics of the queries and passages closely match their training distribution. This suggests that in general-domain texts like Wikipedia, which are linguistically standard and factually dense, the semantic signal of pre-trained models is robust enough, and the additional entity enforcement does not provide the disambiguation benefit observed in the lecture corpus.
A marginal trade-off is observed between the Cross-Encoder variants: while the RRF + Cross-Encoder pipeline achieves a slight advantage in precision metrics, the Standalone version retains a higher Recall@10 ($0.938$). This indicates that RRF acts as a pre-filter, refining the candidate pool for precision at the cost of slightly reducing deep recall.
@@
This divergence provides strong experimental evidence for the Domain Mismatch hypothesis. The Cross-Encoder leverages its pre-training on general web data to rank standard text effectively. However, as demonstrated in the previous sections, this advantage disappears in the high-ambiguity domain of university lectures. This leads to a crucial conclusion: while Cross-Encoders are optimal for general tasks, they are not a universal solution. For specialized educational platforms where training data is scarce, our entity-aware approach provides a more robust retrieval mechanism without the computational cost of a heavy neural re-ranker.
4.5 Discussion
@@
The experimental results present a clear and divergent pattern between the standard SQuAD-it benchmark and the specific educational dataset, which forms a major finding of this study.
First, the quantitative analysis on our specialized educational corpus demonstrates that our proposed ELERAG architecture achieved the best overall performance. It consistently outperformed not only the Weighted-Score baseline but also the Standalone Cross-Encoder. Crucially, while the Standalone Cross-Encoder showed high retrieval capacity (matching ELERAG in Recall), it failed to achieve comparable precision in Exact Match and MRR. This indicates that in a specialized domain with subtle terminological nuances, a generic deep-learning re-ranker may struggle to prioritize the specific "gold" chunk among other semantically plausible candidates. ELERAG, by leveraging the explicit entity signal, successfully bridges this gap.
@@
Conversely, on the general-purpose SQuAD-it benchmark, this trend inverted sharply: the Cross-Encoder configurations obtained the highest scores across all metrics, proving their superiority in a standard, general-domain QA setting. This quantitative divergence provides strong experimental evidence for our Domain Mismatch hypothesis. It indicates that while large, pre-trained re-rankers excel on web-style data (like Wikipedia), a domain-adapted hybrid model like ELERAG is significantly more effective on specialized, narrative corpora (like university lectures). In these contexts, the explicit signals provided by Entity Linking align better with the data than the generic semantic patterns learned by the Cross-Encoder.
@@
Given that ELERAG appears to be the most effective solution for the specific domain, qualitative analysis helps to explain why the integration of Entity Linking provides this advantage. Inspection of the results confirms that the entity-based signal is crucial in high-ambiguity scenarios. A clear example is the query “Who is Smith?”. The Baseline system, relying only on semantic similarity, retrieved scattered and less relevant documents based on broad keyword matching. In contrast, the ELERAG pipeline, guided by the unambiguous entity ID from Wikidata, successfully promoted the correct, relevant passages to the top ranks. This allowed the LLM to generate a more coherent, factually dense, and precise answer. This qualitative benefit was less pronounced in broad, low-ambiguity queries, where the baseline’s semantic search was already sufficient. This suggests that the primary value of our hybrid approach lies in its ability to resolve factual ambiguity, which is a critical weakness of purely semantic systems in technical domains. Moreover, because the system relies on multilingual embeddings and Wikidata identifiers shared across languages, the overall architecture is largely language-agnostic and could be applied to other languages with comparable knowledge base coverage.
Beyond pure retrieval performance, it is crucial to highlight the significant difference in computational cost at inference time. The Standalone Cross-Encoder relies on pairwise evaluations: to re-rank $N$ retrieved chunks, it requires $N$ distinct forward passes through a deep Transformer model analyzing the concatenated query-document strings. Due to the $\mathcal{O}(L^{2})$ complexity of the self-attention mechanism (where $L$ is the sequence token length), this process is computationally expensive and typically requires dedicated GPU hardware to achieve low latency in real-time RAG applications. Conversely, ELERAG shifts the heavy computational burden to the offline indexing phase, where chunk entities are pre-extracted. At query time, the system only processes the short user query through a lightweight NER module, performs standard I/O API calls to Wikidata for the extracted mentions, and computes the final ranking via a simple Boolean set intersection of Wikidata IDs. This highly efficient, CPU-friendly operation avoids heavy neural inference, making it exceptionally suited for scalable, domain-specific deployments.
Another practical consideration concerns the use of automatically transcribed lecture data. Speech recognition may occasionally introduce spelling variations or incomplete entity mentions. While such cases may reduce the effectiveness of the entity linking signal, the hybrid architecture of ELERAG ensures robustness by combining entity-aware retrieval with dense semantic retrieval. This design prevents isolated transcription errors from significantly degrading the overall system performance.
@@
5 Conclusion
We presented ELERAG, an enhanced RAG architecture that integrates a factual signal derived from entity linking to improve the accuracy of educational question-answering systems in Italian. Our system has the potential to enhance RAG systems in domains that differ from standard RAG benchmarks and training data, such as the educational domain.
Our experimental evaluation offers several important implications for research in retrieval-augmented generation and educational AI. First, the consistent improvements obtained through entity-based re-ranking demonstrate that factual grounding can be enhanced without the need for expensive model retraining, solely through structured post-retrieval refinement. Crucially, our experiments demonstrated that injecting structured semantic knowledge (Wikidata) yields superior top-ranking performance (Precision@1 and MRR) compared to strong traditional exact-match baselines like BM25, proving that semantic disambiguation is highly beneficial in specialized educational domains. Second, the contrast between ELERAG and Cross-Encoder performance emphasizes that high-capacity models do not generalize uniformly across all domains: their success strongly depends on the alignment between pre-training data and the target corpus. This finding suggests that lightweight hybrid methods—such as our entity-enriched retrieval—can be a more efficient and accurate alternative to SOTA re-rankers in domain-specific applications, particularly in low-resource or multilingual settings such as Italian educational materials. Finally, the integration of an entity linking module provides a clear framework for combining symbolic and neural representations, showing how explicit knowledge bases like Wikidata can complement dense embeddings to improve interpretability and precision.
Further research directions include exploring adaptive weighting schemes for the RRF, dynamically adjusting the balance between semantic and entity signals based on the query type (e.g., factual vs. conceptual). Moreover, expanding the evaluation to multilingual data and incorporating human-in-the-loop assessments would provide a richer understanding of the system’s real-world educational value.
@@ Acknowledgements.
GPT-4o Generative Pre-trained Transformer 4.0